# Dissecting Mahalanobis: How Feature Geometry and Normalization Shape OOD Detection

### v1

@@ 6 Conclusion
In this work, we conducted a comprehensive empirical study across diverse image foundation models, datasets, and distance normalization schemes to understand how representation geometry shape Mahalanobis-based OOD detection performance. Our comparative analysis revealed that these detectors are not universally reliable, with significant variance in inherent OOD detection capabilities across different self-supervised models. We demonstrated that this variance correlates strongly with measurable geometric properties of the in-distribution feature space. In particular, the product of Local Intrinsic Dimensionality and within-class scatter slope achieves strong correlations with detection effectiveness. The proposed measure reflects the balance between local manifold complexity and cluster compactness, crucial for effective OOD.
Building on these geometric insights, we introduced radially scaled $\ell_{2}$ Normalization, a transformation that provides direct control over the feature space’s radial geometry. This allows to reshape feature distributions to better satisfy the assumptions of the Mahalanobis distance or to improve out-of-distribution (OOD) separation. We also developed a regression framework capable of predicting the optimal $\beta$ value using only in-distribution data, achieving performance comparable to an oracle. Future work should explore fully model-free methods for selecting the optimal $\beta$.
Appendix A Computation and Intuition for Covariance and Scatter Matrices
Our spectral analysis relies on three core matrices derived from the feature embeddings $\{z_{i},y_{i}\}_{i=1}^{N}$ of a trained model: the overall feature covariance $C$, the within-class scatter $S_{w}$, and the between-class scatter $S_{b}$. All eigenvalues reported in the main text are sorted in descending order.
Feature Covariance.
The global covariance is
$C=\frac{1}{N}\sum_{i=1}^{N}(z_{i}-\mu)(z_{i}-\mu)^{\top},\qquad\mu=\frac{1}{N}\sum_{i=1}^{N}z_{i}.$ (3)
This matrix captures the overall spread of representations in feature space. Large eigenvalues correspond to directions of high variance, indicating axes along which the model representation varies the most across all samples.
Within-Class Scatter.
For $K$ classes with means $\mu_{k}$ and $n_{k}$ samples each,
$S_{w}=\frac{1}{N}\sum_{k=1}^{K}\;\sum_{i:y_{i}=k}(z_{i}-\mu_{k})(z_{i}-\mu_{k})^{\top}.$ (4)
$S_{w}$ measures the average dispersion of features around their class means, capturing intra-class variability. Intuitively, if $S_{w}$ has large eigenvalues, samples of a class are more spread out in feature space; small eigenvalues indicate tight, compact clusters. Notably, $S_{w}$ is equivalent to the tied covariance $\Sigma$ used in Mahalanobis distance (Eq. 1) when all classes share a common covariance estimate: $\Sigma=S_{w}$. Thus the Mahalanobis detector implicitly measures distances with respect to the within-class scatter of the training distribution.
Between-Class Scatter.
The between-class scatter quantifies the variability of class means:
$S_{b}=\frac{1}{N}\sum_{k=1}^{K}n_{k}(\mu_{k}-\mu)(\mu_{k}-\mu)^{\top}.$ (5)
This matrix captures inter-class separation: directions with large eigenvalues indicate axes along which class centroids are widely separated, while small eigenvalues correspond to directions where classes overlap. The spectrum of $S_{b}$ provides insight into the model’s ability to linearly discriminate between classes.
@@ Appendix A Computation and Intuition for Covariance and Scatter Matrices
Appendix B Detailed Description of Spectral and Manifold Metrics
Let $X\in\mathbb{R}^{N\times d}$ be the feature matrix and $C=\tfrac{1}{N}(X-\bar{X})^{\top}(X-\bar{X})$ its covariance. All metrics below operate on the sorted non-negative eigenvalues of $C$, denoted $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{d}$.
Intrinsic Dimensionality (ID).
Global estimate of the manifold dimension using maximum-likelihood methods of Ma et al. (2018).
Local Intrinsic Dimensionality (LID).
For a point $z$,
$\operatorname{LID}_{k}(z)=-\!\left[\frac{1}{k}\sum_{j=1}^{k}\log\frac{r_{j}(z)}{r_{k}(z)}\right]^{-1},$
where $r_{j}$ is the $j$-th nearest-neighbor distance. We report the dataset mean for $k\!\in\!\{10,25,50,100\}$.
Total Variance.
$\text{Total Variance}=\sum_{i=1}^{d}\lambda_{i}.$
Effective Rank.
Implements the code’s definition,
$\text{Effective Rank}=\frac{\sum_{i}\lambda_{i}}{\lambda_{1}},$
the ratio of total variance to the largest eigenvalue.
Participation Ratio.
$\text{PR}=\frac{\bigl(\sum_{i}\lambda_{i}\bigr)^{2}}{\sum_{i}\lambda_{i}^{2}}.$
Condition Number.
$\kappa=\frac{\lambda_{1}}{\lambda_{d}},$
where $\lambda_{d}$ is the smallest positive eigenvalue.
Spectral Gap.
Difference between the largest and sixth largest eigenvalue,
$\text{Gap}=\lambda_{1}-\lambda_{6}.$
Entropy.
Shannon entropy of the normalized spectrum,
$\text{Entropy}=-\sum_{i}p_{i}\log p_{i},\qquad p_{i}=\frac{\lambda_{i}}{\sum_{j}\lambda_{j}}.$
Average Log Decay Rate (Top-20).
Mean forward difference of the first 20 log-eigenvalues,
$\frac{1}{19}\sum_{i=1}^{19}\bigl[\log\lambda_{i}-\log\lambda_{i+1}\bigr].$
Slope.
Slope of the least-squares fit
$\log\lambda_{i}=a+b\,i,$
i.e., regression of $\log\lambda_{i}$ on the linear index $i$.
Beta Power Law.
Exponent $\beta$ of a power law $\lambda_{i}\propto i^{-\beta}$, computed as the negative slope of $\log\lambda_{i}$ versus $\log i$.
Dim. 90% Var.
Minimum $k$ such that $\sum_{i=1}^{k}\lambda_{i}\big/\sum_{j}\lambda_{j}\geq 0.9$.
These definitions ensure exact reproducibility of the results reported in Section 4.3.
@@ Appendix B Detailed Description of Spectral and Manifold Metrics
Appendix C Geometry Induced by Radially Scaled $\ell_{2}$ Normalization
This appendix provides a formal derivation of the Riemannian metric induced by the tunable radial $\ell_{2}$ normalization used in our Mahalanobis-based OOD detector.
C.1 Setup and Derivation

### v3

@@ 8 Conclusion
We studied Mahalanobis-style OOD detection across a range of vision foundation-model backbones, OOD benchmarks, and feature normalizations, and found that performance depends strongly on the representation. Across models and detector variants, two ID geometry signals, local intrinsic dimensionality and the within-class spectral slope, track this variation. We also introduced radially scaled $\ell_{2}$ normalization, which adjusts feature radii while keeping directions fixed, and an ID-only rule for choosing $\beta$. This selection improves over fixed normalizations in most settings and can approach oracle-tuned performance without using OOD samples.
Impact Statement
This paper studies how representation geometry and feature normalization affect Mahalanobis-style OOD detection, with the goal of improving the reliability of deployed vision models. The primary positive impact is practical: our findings provide diagnostics and simple post-hoc normalization procedures that can reduce false positives and improve robustness monitoring in safety-relevant settings (e.g., medical imaging, autonomous systems, industrial inspection). Potential risks include misuse for overconfidence: improved OOD detection may be treated as a guarantee of safety, despite the fact that OOD detection remains imperfect and depends on data, model, and deployment conditions. In addition, geometry-based tuning could be adapted to evade certain detectors if an adversary can influence feature distributions. We therefore emphasize that our methods should be used as one component of a broader reliability pipeline (e.g., calibration, auditing, and monitoring), and that deployment decisions should not rely on OOD scores alone.
Appendix A Detailed Description of Spectral and Manifold Metrics
All spectral metrics are computed from the ordered eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{d}$ of a chosen matrix $M\in\{C,S_{w},S_{b}\}$ (see Appendix B for definitions). When needed, we use the normalized spectrum $p_{i}\triangleq\lambda_{i}/\sum_{j=1}^{d}\lambda_{j}$.
Intrinsic dimensionality (ID).
A global estimate of manifold dimension using the maximum-likelihood estimator of Ma et al. (2018).
Local intrinsic dimensionality (LID).
For a feature vector $z$, let $r_{j}(z)$ denote the distance to its $j$-th nearest neighbor in feature space. The kkNN LID estimator is
$\mathrm{LID}_{k}(z)\;=\;-\left[\frac{1}{k}\sum_{j=1}^{k}\log\frac{r_{j}(z)}{r_{k}(z)}\right]^{-1}.$ (16)
We report the dataset mean for $k\in\{10,25,50,100\}$.
Total variance (trace).
$\mathrm{Tr}(M)\;=\;\sum_{i=1}^{d}\lambda_{i}.$ (17)
Effective rank (trace-to-top ratio).
We use the ratio
$r_{\mathrm{eff}}\;\triangleq\;\frac{\sum_{i=1}^{d}\lambda_{i}}{\lambda_{1}}.$ (18)
Participation ratio (PR).
$\mathrm{PR}\;\triangleq\;\frac{\left(\sum_{i=1}^{d}\lambda_{i}\right)^{2}}{\sum_{i=1}^{d}\lambda_{i}^{2}}.$ (19)
Condition number.
$\kappa\;\triangleq\;\frac{\lambda_{1}}{\lambda_{d}},$ (20)
where $\lambda_{d}$ is the smallest (non-zero) eigenvalue (or the smallest eigenvalue after numerical regularization, when applicable).
Spectral gap (head).
$\mathrm{Gap}\;\triangleq\;\lambda_{1}-\lambda_{6}.$ (21)
Spectral entropy.
$\mathrm{H}\;\triangleq\;-\sum_{i=1}^{d}p_{i}\log p_{i},\qquad p_{i}=\frac{\lambda_{i}}{\sum_{j}\lambda_{j}}.$ (22)
Average log decay rate (top-20).
$\frac{1}{19}\sum_{i=1}^{19}\bigl(\log\lambda_{i}-\log\lambda_{i+1}\bigr).$ (23)
Log-spectrum slope (full-range).
We fit a least-squares line to the log-spectrum
$\log\lambda_{i}\;=\;a+b\,i,$ (24)
and report the fitted slope $b$ (in the main paper we often use $|s|$ for the magnitude of this slope when $s<0$).
Power-law exponent.
We fit $\lambda_{i}\propto i^{-\beta}$ by regressing $\log\lambda_{i}$ on $\log i$ and report the exponent $\beta$.
Dimension for 90% explained variance.
$k_{0.9}\;\triangleq\;\min\left\{k:\frac{\sum_{i=1}^{k}\lambda_{i}}{\sum_{j=1}^{d}\lambda_{j}}\geq 0.9\right\}.$ (25)
These definitions match the metrics used in Section 5.2 and enable exact reproducibility.
Appendix B Computation and Intuition for Covariance and Scatter Matrices
Our spectral analyses are based on three symmetric positive semidefinite matrices computed from in-distribution (ID) feature embeddings. Let $z_{i}\in\mathbb{R}^{d}$ be the feature of sample $x_{i}$ with label $y_{i}\in\{1,\dots,K\}$, and let $N$ be the number of ID samples. Define the global mean $\mu\triangleq\tfrac{1}{N}\sum_{i=1}^{N}z_{i}$, class means $\mu_{k}\triangleq\tfrac{1}{n_{k}}\sum_{i:y_{i}=k}z_{i}$, and class counts $n_{k}$. All eigenvalues are reported in descending order, $\lambda_{1}\geq\cdots\geq\lambda_{d}$.
Global covariance ($C$).
We measure the overall spread of ID features with
$C\;\triangleq\;\frac{1}{N}\sum_{i=1}^{N}(z_{i}-\mu)(z_{i}-\mu)^{\top}.$ (26)
Large eigenvalues of $C$ correspond to directions of high variance across the entire ID dataset.
Within-class scatter ($S_{w}$).
To measure intra-class variability we use the tied within-class scatter
$S_{w}\;\triangleq\;\frac{1}{N}\sum_{k=1}^{K}\;\sum_{i:y_{i}=k}(z_{i}-\mu_{k})(z_{i}-\mu_{k})^{\top}.$ (27)
When the Mahalanobis detector is implemented with a shared (tied) covariance across classes, its covariance estimate coincides with $S_{w}$. In particular, the standard class-conditional MD score can be written as $\mathrm{MD}_{k}(z)=(z-\mu_{k})^{\top}S_{w}^{-1}(z-\mu_{k})$, matching Eq. 3 in the main text.
Between-class scatter ($S_{b}$).
To quantify how class means spread around the global mean, we use
$S_{b}\;\triangleq\;\frac{1}{N}\sum_{k=1}^{K}n_{k}\,(\mu_{k}-\mu)(\mu_{k}-\mu)^{\top}.$ (28)
Large eigenvalues of $S_{b}$ indicate directions along which class centroids are well separated.
@@ Appendix B Computation and Intuition for Covariance and Scatter Matrices
Appendix C Spectral analysis of training effects
To understand how the intrinsic geometry of representations affects OOD performance, we begin by examining the spectral properties of three key matrices: the feature covariance $C$, the within-class scatter $S_{w}$, and the between-class scatter $S_{b}$. These matrices capture complementary aspects of the feature space: $C$ reflects overall variance, $S_{w}$ measures intra-class dispersion, and $S_{b}$ quantifies inter-class separation (more details in Appendix B). Our first analysis focuses on the eigenvalue spectra of these matrices. The magnitude and decay of eigenvalues reveal how variance is distributed across dimensions, providing insight into the richness and anisotropy of the feature space. For instance, a steep decay in $S_{w}$ eigenvalues indicates that intra-class variability is concentrated along a few directions, resulting in tight clusters, whereas a slower decay suggests more diffuse intra-class variation. Similarly, large eigenvalues in $S_{b}$ correspond to well-separated class means, signaling strong discriminability.
Spectral ratios
Eigenvalue shifts
Beyond static spectra, we are interested in how stable the representation geometry is under distributional shifts. To capture this, we define a spectral shift metric, which measures the relative change in eigenvalues from the training set to validation or OOD data (see Appendix B). A small shift indicates that the representation preserves its structure across data splits, signaling robustness. Large positive shifts reveal that features are spreading along new directions, while large negative shifts indicate compression. Figure 10 shows that OOD samples induce larger spectral shifts in models trained on small datasets, reflecting lower generalization and brittle feature structures. Large-scale pretrained models show smaller shifts, indicating more stable, robust representations under distributional change. Fine-tuning generally maintains small shifts while increasing $S_{b}$, improving class separation without compromising intra-class compactness.
@@ Appendix C Spectral analysis of training effects
Appendix D Geometry Induced by Radially Scaled $\ell_{2}$ Normalization and $\beta$ Selection
This appendix provides geometric intuition for the radially scaled $\ell_{2}$ map $\phi_{\beta}$ (Sec. 6.2) and specifies the ID-only $\beta$ selection rule used throughout the paper (Sec. 6.4). Our goal is modest: (i) show that $\beta$ continuously reweights radial versus angular variations, explaining why both LID $m(\beta)$ and within-class spectral summaries $s(\beta)$ vary smoothly with $\beta$; and (ii) formalize how we choose $\hat{\beta}$ from the proxy curve $P(\beta)$ without committing a priori to “minimize” or “maximize.”
D.1 Setup and induced metric